(a) Let z1 = 1 + √2i and 1 −√2i.
(i) Determine the polar form of z1.
(ii) Determine that the polar form of z2.
(iii) Use the polar forms of z1 and z2 to verify that z1 · z2 = 3
(iv) Use the polar forms of z1 and z2 to verify that −1/3 +2/3√2i =z1/z2
1
Expert's answer
2020-10-14T14:24:00-0400

"\displaystyle(a) \\

z_1 = 1 + i\sqrt{2}, z_2 = 1 - i\sqrt{2}\\


(i) \\
\textsf{Polar form of}\, z_1 \\

\begin{aligned}
r &= \sqrt{1 + (\sqrt{2})^2}\\
&=\sqrt{1 + 2} \\
&= \sqrt{3}
\end{aligned} \\

\begin{aligned}
\theta = \arctan\left(\frac{\sqrt{2}}{1}\right) &= \arctan(\sqrt{2})\\
&= 54.74\degree
\end{aligned} \\

\therefore z_1 = \sqrt{3}(\cos(54.74\degree) + i\sin(54.74\degree)) \\

(ii)\\
\textsf{Polar form of}\, z_2\\

\begin{aligned}
r &= \sqrt{1 + (\sqrt{2})^2}\\
&=\sqrt{1 + 2} \\
&= \sqrt{3}
\end{aligned} \\

\textsf{Since}\, \theta\,\textsf{is in the fourth quadrant,}\\ 

\begin{aligned}
\theta &=360\degree - \arctan\left(\frac{\sqrt{2}}{1}\right) 
\\&= 360\degree - \arctan(\sqrt{2}) 
\\&= 360\degree - 54.74\degree
\\&= 305.26\degree
\end{aligned}\\

\therefore z_2 = \sqrt{3}(\cos(305.26\degree) + i\sin(305.26\degree)) \\

(iii)\\\begin{aligned}
z_1 \cdot z_2 &= \sqrt{3}(\cos(54.74\degree) + i\sin(54.74\degree)) \\&\times \sqrt{3}(\cos(305.26\degree) + i\sin(305.26\degree))
\\&= 3(\cos(54.74\degree)\cos(305.26\degree) \\&+ i(\cos(54.74\degree)\sin(305.26\degree) + \sin(54.74\degree)\cos(305.26\degree)) \\&- \sin(54.74\degree)\sin(305.26\degree))
\end{aligned}\\

\begin{aligned}
\cos(305.26\degree) &= \cos(360\degree - 54.74\degree) = \cos(54.74\degree) \\
\sin(305.26\degree) &= -\sin(360\degree - 54.74\degree) = -\sin(54.74\degree) 
\end{aligned}\\


\begin{aligned}
z_1 \cdot z_2 &= 3(\cos(54.74\degree)\cos(54.74\degree) \\&+ i(\cos(54.74\degree)\cdot -\sin(54.74\degree) + \sin(54.74\degree)\cos(54.74\degree)) \\&- \sin(54.74\degree)\cdot-\sin(54.74\degree))
\\&= 3(\cos^2(54.74\degree) + \sin^2(54.74\degree)) = 3\times 1 = 3
\end{aligned}\\

(iv) \\
\begin{aligned}
\frac{z_1}{z_2} &= \frac{\sqrt{3}(\cos(54.74\degree) + i\sin(54.74\degree))}{\sqrt{3}(\cos(305.26\degree) + i\sin(305.26\degree))}
\\&= \frac{(\cos(54.74\degree) + i\sin(54.74\degree))}{(\cos(305.26\degree) + i\sin(305.26\degree))} \times \frac{(\cos(305.26\degree) - i\sin(305.26\degree))}{(\cos(305.26\degree) - i\sin(305.26\degree))}
\\&=\frac{\cos(54.74\degree)\cos(305.26\degree)}{{(\cos^2(305.26\degree) + \sin^2(305.26\degree)}}+\\& \frac{i(-\cos(54.74\degree)\sin(305.26\degree) + \sin(54.74\degree)\cos(305.26\degree))}{(\cos^2(305.26\degree) + \sin^2(305.26\degree)} \\&+ \frac{\sin(54.74\degree)\sin(305.26\degree)}{(\cos^2(305.26\degree) + \sin^2(305.26\degree)}
\\&=\frac{\cos^2(54.74\degree) + i(\cos(54.74\degree)\sin(54.74\degree))}{1} 
\\&+ \frac{\sin(54.74\degree)\cos(54.74\degree)) - \sin^2(54.74\degree)}{1} 
\\&= \cos^2(54.74\degree) - \sin^2(54.74\degree) + 2i(\cos(54.74\degree)\sin(54.74\degree)) 
\\&= \cos^2(\arctan(\sqrt{2})) - \sin^2(\arctan(\sqrt{2})) \\&\hspace{4.5cm}+ 2i(\cos(\arctan(\sqrt{2}))\sin(\arctan(\sqrt{2})))
\end{aligned}\\


\textsf{Let}\, y = \cos(\arctan(\sqrt{x})), u = \arctan(\sqrt{x}) \\

y = \cos{u}, x = \tan{u} \\

1 + \tan^2{u} = \sec^2{u} \implies \cos{u} = \frac{1}{\sqrt{1 + \tan^2{u}}}\\

\therefore y = \cos(\arctan(\sqrt{x})) = \frac{1}{\sqrt{1 + \tan^2{u}}}= \frac{1}{\sqrt{1 + x^2}}\\

\therefore \cos^2(\arctan(\sqrt{2})) = \frac{1}{1 + (\sqrt{2})^2} = \frac{1}{3} \, \& \\
\sin^2(\arctan(\sqrt{2})) = 1 - \cos^2(\arctan(\sqrt{2})) = 1 - \frac{1}{3} = \frac{2}{3} \\

\begin{aligned}
\therefore \frac{z_1}{z_2} &= \frac{1}{3} - \frac{2}{3} + 2i\left(\sqrt{\frac{2}{3}\cdot\frac{1}{3}}\right)
\\&= -\frac{1}{3} + i\frac{2}{3}\sqrt{2}
\end{aligned}"